OpenStudy (anonymous):

.. What is the simplest form of the number? 16^3/4

OpenStudy (wolf1728):

16^3/4 =16*16*16 /4 =16*16*4 =256*4 1,024 = 2^10

OpenStudy (bibby):

I'm pretty sure it's 8. it's \[\sqrt[4]{16^3} = \sqrt{\sqrt{4096}}\]